Please solve : log(2) (x+5) - log(2) (2x)=8
Answer by ewatrrr(24785)

log(2) (x+5) - log(2) (2x)=8 Note: 2^8 = 256
(x+5)/2x = 256
x+5 = 512x
5/511 = x
.......
***


***